# allinspirekeys.bib

### v1

@@ Abstract
At finite lattice spacing, Lagrangian and Hamiltonian predictions differ due to discretization effects. In the Hamiltonian limit, i.e. at vanishing temporal lattice spacing $a_{t}$, the path integral approach in the Lagrangian formalism reproduces the results of the Hamiltonian theory. In this work, we numerically calculate the Hamiltonian limit of a U$(1)$ gauge theory in $(2+1)$ dimensions. This is achieved by Monte Carlo simulations in the Lagrangian formalism with lattices that are anisotropic in the time direction. For each ensemble, we determine the ratio between the temporal and spatial scale with the static quark potential and extrapolate to $a_{t}\to 0$. Our results are compared with the data from Hamiltonian simulations at small volumes, showing agreement within $<2\sigma$. These results can be used to match the two formalisms.
@@ 1 Introduction
Gauge theories are fundamental in our understanding of force mediation in the standard model (SM) of particle physics. Of the three forces unified in the SM, the strong force or quantum chromodynamics (QCD) is special because it is strongly coupled in the low energy regime. Therefore, it requires a non-perturbative treatment, which is possible in the lattice regularisation of gauge theories pioneered by Wilson [Wilson:1974sk]. While primarily applied to QCD, the lattice regularisation is applicable to any gauge theory.
@@
In practice, any computation in a lattice gauge theory requires one to choose either the path integral formalism enabling mainly Monte Carlo (MC) simulations of such theories, or the Hamiltonian formalism. The basis for the former has been provided already by Wilson, for the latter the corresponding Hamiltonian has been derived not much later based on general arguments in Ref. [Kogut:1974ag] by Kogut and Susskind, while Creutz derived the same expression for the Hamiltonian by explicitly constructing the transfer matrix [Creutz:1976ch]. The Hamiltonian formulation has recently attracted fresh attention since it represents the natural formulation one would use on a future digital quantum computer. Compared to the MC approach, Hamiltonian simulations have the advantage that for instance systems at finite density or real time evolution can be studied. However, the development state of current quantum computing devices limits such simulations to rather small systems. Alternatively, tensor network states can be used, which see rapid development as well [Banuls:2018jag, Banuls:2019rao, Magnifico:2020bqt, Felser:2019xyv, Magnifico:2024eiy]. Still, Hamiltonian simulations are currently restricted to systems with a relatively small number of degrees of freedom.
@@
Ideas to nevertheless usefully apply Hamiltonian simulations already now include a clever combination with MC simulations, profiting from the respective strengths simultaneously [Clemente:2022cka, Crippa:2024cqr, Avkhadiev:2019niu, Avkhadiev:2022ttx, Avkhadiev:2024yjs]. One such idea has been brought forward in Ref. [Clemente:2022cka] and further investigated in Ref. [Crippa:2024cqr]. It requires the matching of Hamiltonian and Lagrangian simulations: both formulations encompass the bare gauge coupling as a single parameter, which is directly related to the scale of the corresponding theory. However, being a bare parameter implies that simply using the same numerical value for the coupling will, at least at finite lattice spacing, likely lead to sizeable artefacts. An alternative is to connect the MC simulations with the Hamiltonian ones by taking the continuum limit in time direction, as the construction by Creutz suggests. This continuum limit in time direction has been studied previously in Refs. [Loan:2002ej, Byrnes:2003gg, Loan:2003wy] based on the so-called anisotropic formulation of lattice gauge theories [Morningstar:1999rf]. Most relevant for our work here is Ref. [Loan:2002ej], where the authors study a U$(1)$ theory using the anisotropic Wilson plaquette action. They take the temporal continuum limit keeping the $\beta$-value fixed. They compare to Green’s Function Monte Carlo results which they comment, however, to be unreliable due to strong dependence on the trial wave function.
@@
In this paper we will go beyond Refs. [Loan:2002ej, Byrnes:2003gg] in two ways: first, we will take the continuum limit in time direction in compact pure U$(1)$ gauge theory while keeping a suitable spatial length fixed, which we determine non-perturbatively. Second, we directly compare our extrapolated results from MC simulations with Hamiltonian simulations, finding agreement within $2\sigma$. It is conceptually straightforward to extend this to non-Abelian lattice gauge theories. We reported on a first stage of this work in Ref. [Funcke:2022opx].
@@ 2.1 Lagrangian formulation
On the Lagrangian side, we use the anisotropic Wilson action [Morningstar:1997ff, Morningstar:1999rf], which reduces to the standard Wilson action for the special value of the anisotropy parameter $\xi_{\text{input}}=1$. It reads:
@@
$\xi_{\text{input}}$ is the bare anisotropy and $\beta=1/g_{0}^{2}$ the inverse squared coupling constant. The gauge links $U_{\mu}(\vec{x})$ are elements of U$(1)$ and can, hence, be parametrised as $U=e^{i\varphi}$ with a real-valued angle $\varphi$. $\vec{x}$ is a point in our $2+1$ dimensional lattice and the directions $\mu\in\{0,1,2\}$. MC simulations of the theory described by the action $S_{W}$ can be performed using standard Markov Chain MC methods, such as the Metropolis algorithm. More details on the algorithm will be given below.
@@
In this paper we use two different choices for such an observable $O$, both based on the so-called static quark potential $V$, see for instance Ref. [Loan:2002ej]. The static quark potential can be determined from planar Wilson loops $W(a_{\mu}x,a_{\nu}y)$ with extents $a_{\mu}x$ and $a_{\nu}y$. Here, $a_{\mu}$ represents the lattice spacing in direction $\mu$, which can be one of the spatial directions or the time direction. Since we are working in Euclidean space-time, the expectation values of Wilson loops decay exponentially in spatial as well as temporal directions. By forming purely spatial Wilson loops $W_{ss}$ and spatial-temporal Wilson loops $W_{st}$, one obtains
$\begin{split}\lim_{y\to\infty}\frac{W_{ss}(x/a_{s},(y+1)/a_{s})}{W_{ss}(x/a_{s},y/a_{s}y)}&=\exp(-a_{s}V_{s}(x/a_{s}))\,,\\ \lim_{t\to\infty}\frac{W_{st}(x/a_{s},(t+1)/a_{t})}{W_{st}(x/a_{s},t/a_{t})}&=\exp(-a_{t}V_{t}(x/a_{s}))\,.\end{split}$ (4)
@@
to the data for the two potentials as a function of distance. The fit parameter $c$ represents the difference in self-energy in $V_{s}$ and $V_{t}$. We refer to this procedure as the one based on the “normal” potential.
The second way to determine the potential is the one also used by Ref. [Alford:2000an], and we refer to this method as the one based on the “sideways” potential. In this procedure the potential is determined varying the first argument in the corresponding Wilson loops as follows
@@ 2.1.1 Sommer parameter $r_{0}$ and setting the scale
can be used to define a length scale $r_{0}$, the so-called Sommer parameter [Sommer:1993ce], as follows
$-r^{2}\frac{\mathrm{d}}{\mathrm{d}r}V(r)|_{r=r_{0}}=c$ (9)
in units of the spatial lattice spacing $a_{s}$. In QCD, the physical value of $r_{0}$ is known to be around $0.5\ \mathrm{fm}$ for $c=-1.65$, but, in the U$(1)$ theory at hand its physical value is unknown. However, this is not relevant for our procedure, as we only need an observable in units of the spatial lattice spacing with a well defined continuum limit. We use a value of $c=-1.65$, because it turns out to be in the linear region of the potential and we stick to the notation $r_{0}/a_{s}$. Once the potentials are parametrised using the form eq. 8, the corresponding value of $r_{0}/a_{s}$ can be determined. It allows one to fix the spatial lattice spacing: if in two simulations with parameters $(\beta,\xi)$ and $(\beta^{\prime},\xi^{\prime})$ the same $r_{0}/a_{s}$ is measured within uncertainties, both simulations exhibit the same $a_{s}$.
@@ 2.2 Hamiltonian
The Kogut-Susskind Hamiltonian for the pure U$(1)$ lattice gauge theory in $(2+1)$ dimensions is given by [Kogut:1974ag]
@@
where $\vec{A}_{\vec{r},\mu}$ is the discretized vector field in the compact formulation, i.e., the values of $ag\vec{A}_{\vec{r},\mu}$ are constrained to lie within the interval $[0,2\pi)$. Note that $\hat{U}_{\vec{r},\mu}$ is a unitary operator. The commutation relations between the electric field operator $\hat{E}_{\vec{r},\nu}$ and the link operator $\hat{U}_{\vec{r^{\prime}},\mu}$ read
@@
$\Bigg{[}\sum\limits_{\mu=1,2}\left(\hat{E}_{\vec{r},\mu}-\hat{E}_{\vec{r}-\mu,\mu}\right)$ $-Q_{\vec{r}}\Bigg{]}\ket{\Phi}=0,$ (15)
where $Q_{\vec{r}}$ are the static charges. Instead of considering the full Hilbert space and enforcing Gauss’s law, we formulate the theory directly on the gauge-invariant subspace, by using the Gauss’s law constraints to eliminate certain degrees of freedom [Haase:2020kaj, Kaplan:2018vnj, Paulson:2020zjd, Kan:2021nyu, Bauer:2021gek].
For a numerical implementation of the Hamiltonian, the gauge degrees of freedom must be truncated to a finite dimension because the electric field values are unbounded, resulting in an infinite-dimensional Hilbert space for these degrees of freedom. The continuous U$(1)$ gauge group can be discretized in the electric basis to the group of integers $\mathbb{Z}_{2l+1}$. The integer $l$ sets the truncation level, i.e., the discretized gauge fields are constrained to integer values within the range $[-l,l]$ [Haase:2020kaj]. The total Hilbert space dimension is $(2l+1)^{N}$, where $N$ is the number of gauge fields.
@@
$\hat{U}_{\vec{r},\mu}\ket{e_{\vec{r},\mu}}=\ket{e_{\vec{r},\mu}+1},\ \ \ \hat{U}^{{\dagger}}_{\vec{r},\mu}\ket{e_{\vec{r},\mu}}=\ket{e_{\vec{r},\mu}-1}.$ (17)
When discretizing a gauge theory, one needs to give up either unitarity or the exact commutation relations between the electric field and link operators in eq. 13 and eq. 14. In our case, the commutation relations are preserved for the truncated operators, but unitarity is lost, $\hat{U}^{{\dagger}}_{\vec{r},\mu}\hat{U}_{\vec{r},\mu}\neq\mathbbm{1}$. This can be seen from the matrix representation of the link operators [Paulson:2020zjd],
@@
However, unitarity is recovered in the limit of $l\to\infty$. The resulting errors due to the finite truncation parameter $l$ have been investigated in Refs. [Kuhn:2014rha, Buyens:2017crb]. Alternative methods for defining the electric field and link operators have been explored in Refs. [Mathis:2020fuo, Chandrasekharan:1996ih, Wiese:2013uua, Hashizume:2021qbb].
In order to analytically derive the Kogut-Susskind Hamiltonian in eq. 11 from the Wilson action in eq. 1, one needs to employ the transfer matrix method [Creutz:1976ch]. This derivation has been performed for various quantum field theories in arbitrary dimensions, for example, for studying transport coefficients [Cohen:2021imf] and the topological $\theta$-term of (non)Abelian lattice gauge theories in (3+1) dimensions [Kan:2021nyu]. Since the Wilson action is defined on a $(d+1)$-dimensional space-time lattice and the Kogut-Susskind Hamiltonian is defined on a $d$-dimensional spatial lattice, the limit of $a_{t}\to 0$ has to be taken when deriving eq. 11 from eq. 1 using the transfer matrix formalism [Creutz:1976ch].
@@ 3.1 Lagrangian
Details of the algorithm can be found in Ref. [Hattori:1992qk] (see in particular the arXiv version). Our implementation can be found in our publicly available simulation code [Urbach_Gross_Romiti]. Our code for the data analysis uses the code [kostrzewa_2024_14800686] and can be found in  [^1] https://github.com/christianegross/U1_analyse_potential.
@@ 3.2 Hamiltonian
As discussed in section 2.2, we use the Gauss’s law to eliminate some gauge degrees of freedom, thereby restricting the theory to the gauge-invariant space. More specifically, we treat the Gauss’s law in eq. 15 as a set of constraints on the electric operators, and solve this set of equations over the electric operators. While there are $N$ Gauss’s law constraints, they are not independent, since there is a conservation of charges, which, in the pure gauge case, means the constraints sum to zero. Therefore, there are only $N-1$ independent constraints, which allows us to express $N-1$ arbitrary electric field operators, i.e., effectively eliminating them, in terms of the remaining ones. Since the eliminated electric fields do not contribute directly to the dynamics, their corresponding link operators become identities. For a two-dimensional $L\times L$ square lattice, where $N=L^{2}$, $L^{2}-1$ out of the $2L^{2}$ gauge degrees of freedom are eliminated. Thus, the Hamiltonian is expressed in terms of the $L^{2}+1$ remaining gauge fields. This reduces the number of basis states from $(2l+1)^{2L^{2}}$ to $(2l+1)^{L^{2}+1}$, which in practice, alleviates the computational costs significantly.
@@
where $V$ is the number of plaquettes in the lattice. Note that we set the lattice spacing $a=1$ throughout our simulations. The simulations are carried out over a range of $1/g^{2}$ in $(0,10]$ and for $l\in\{1,2,3,4\}$. We find that the values of $\langle\mathcal{P}\rangle$ obtained for $l=2,3,4$ agree with each other up to $1.51/g^{2}=$ 1.5, beyond which the values start to deviate for different $l$ values, indicating that the simulations are no longer reliable.

### v2

# Matching Lagrangian and Hamiltonian Simulations in (2+1)-dimensional U(1) Gauge Theory

@@ Abstract
At finite lattice spacing, Lagrangian and Hamiltonian predictions differ due to discretization effects. In the Hamiltonian limit, i.e. at vanishing temporal lattice spacing $a_{t}$, the path integral approach in the Lagrangian formalism reproduces the results of the Hamiltonian theory. In this work, we numerically calculate the Hamiltonian limit of a $U(1)$ gauge theory in $(2+1)$ dimensions. This is achieved by Monte Carlo simulations in the Lagrangian formalism with lattices that are anisotropic in the time direction. For each ensemble, we determine the ratio between the temporal and spatial scale with the static quark potential and extrapolate to $a_{t}\to 0$. Our results are compared with the data from Hamiltonian simulations at small volumes, showing agreement within $2\sigma$. These results can be used to match the two formalisms.
@@ 1 Introduction
Gauge theories are fundamental in our understanding of force mediation in the standard model (SM) of particle physics. Of the three forces unified in the SM, the strong force or quantum chromodynamics (QCD) is special because it is strongly coupled in the low energy regime. Therefore, it requires a non-perturbative treatment, which is possible in the lattice regularisation of gauge theories pioneered by Wilson [1]. While primarily applied to QCD, the lattice regularisation is applicable to any gauge theory.
@@
In practice, any computation in a lattice gauge theory requires one to choose either the path integral formalism enabling mainly Monte Carlo (MC) simulations of such theories, or the Hamiltonian formalism. The basis for the former has been provided already by Wilson, for the latter the corresponding Hamiltonian has been derived not much later based on general arguments in Ref. [2] by Kogut and Susskind, while Creutz derived the same expression for the Hamiltonian by explicitly constructing the transfer matrix [3]. The Hamiltonian formulation has recently attracted fresh attention since it represents the natural formulation one would use on a future digital quantum computer. Compared to the MC approach, Hamiltonian simulations have the advantage that for instance systems at finite density or real time evolution can be studied. However, the development state of current quantum computing devices limits such simulations to rather small systems. Alternatively, tensor network states can be used, which see rapid development as well [4, 5, 6, 7, 8]. Still, Hamiltonian simulations are currently restricted to systems with a relatively small number of degrees of freedom.
@@
Ideas to nevertheless usefully apply Hamiltonian simulations already now include a clever combination with MC simulations, profiting from the respective strengths simultaneously [9, 10, 11, 12, 13, 14, 15]. One such idea has been brought forward in Ref. [9] and further investigated in Ref. [10]. It requires the matching of Hamiltonian and Lagrangian simulations: both formulations encompass the bare gauge coupling as a single parameter, which is directly related to the scale of the corresponding theory. However, being a bare parameter implies that simply using the same numerical value for the coupling will, at least at finite lattice spacing, likely lead to sizeable artefacts. An alternative is to connect the MC simulations with the Hamiltonian ones by taking the continuum limit in time direction, as the construction by Creutz suggests. This continuum limit in time direction has been studied previously in Refs. [16, 17, 18] based on the so-called anisotropic formulation of lattice gauge theories [19]. Most relevant for our work here is Ref. [16], where the authors study a U(1)(1) theory using the anisotropic Wilson plaquette action. They take the temporal continuum limit keeping the $\beta$-value fixed. They compare to Green’s Function Monte Carlo results which they comment, however, to be unreliable due to strong dependence on the trial wave function.
@@
In this paper we will go beyond Refs. [16, 17] in two ways: first, we will take the continuum limit in time direction in compact pure U(1)(1) gauge theory while keeping a suitable spatial length fixed, which we determine non-perturbatively. Second, we directly compare our extrapolated results from MC simulations with Hamiltonian simulations, finding agreement within $2\sigma$. It is conceptually straightforward to extend this to non-Abelian lattice gauge theories. The continuum limit $a_{s}\to 0$ is not considered in this work, we rather want to compare the two formulations at fixed $a_{s}$. We reported on a first stage of this work in Ref. [20].
@@ 2.1 Lagrangian formulation
On the Lagrangian side, we use the anisotropic Wilson action [21, 19], which reduces to the standard Wilson action for the special value of the anisotropy parameter $\xi_{\text{input}}=1$. It reads:
@@
$\xi_{\text{input}}$ is the bare anisotropy and $\beta=1/g_{0}^{2}$ the inverse squared coupling constant. The gauge links $U_{\mu}(\vec{x})$ are elements of U(1)(1) and can, hence, be parametrised as $U=e^{i\varphi}$ with a real-valued angle $\varphi$. $\vec{x}$ is a point in our $2+1$ dimensional lattice and the directions $\mu\in\{0,1,2\}$. MC simulations of the theory described by the action $S_{W}$ can be performed using standard Markov Chain MC methods, such as the Metropolis algorithm. More details on the algorithm will be given below.
@@
In this paper we use two different choices for such an observable $O$, both based on the so-called static quark potential $V$, see for instance Ref. [16]. The static quark potential can be determined from planar Wilson loops $W(a_{\mu}x,a_{\nu}y)$ with extents $a_{\mu}x$ and $a_{\nu}y$. Here, $a_{\mu}$ represents the lattice spacing in direction $\mu$, which can be one of the spatial directions or the time direction. Since we are working in Euclidean space-time, the expectation values of Wilson loops decay exponentially in spatial as well as temporal directions. By forming purely spatial Wilson loops $W_{ss}$ and spatial-temporal Wilson loops $W_{st}$, one obtains
$\begin{split}\lim_{y\to\infty}\frac{W_{ss}(x/a_{s},(y+1)/a_{s})}{W_{ss}(x/a_{s},y/a_{s})}&=\exp(-a_{s}V_{s}(x/a_{s}))\,,\\ \lim_{t\to\infty}\frac{W_{st}(x/a_{s},(t+1)/a_{t})}{W_{st}(x/a_{s},t/a_{t})}&=\exp(-a_{t}V_{t}(x/a_{s}))\,.\end{split}$ (4)
@@
to the data for the two potentials as a function of distance. The fit parameter $c$ represents the difference in self-energy in $V_{s}$ and $V_{t}$ and arises from the fact that the potentials are measured along different axes. The self energy depends on the direction in which the quarks propagate [22]. In the above determination, the quarks propagate in the temporal direction in the measurement of the temporal potential, and propagate in the spatial direction in the measurement of the spatial potential. We refer to this procedure as the one based on the “normal” potential.
The second way to determine the potential is the one also used by Ref. [23], and we refer to this method as the one based on the “sideways” potential. In this procedure the potential is determined varying the first argument in the corresponding Wilson loops as follows
@@
Using the determinations of the Wilson loop above, the quarks propagate in the spatial direction in the measurement of both the spatial and temporal potential. Therefore the potentials have the same self energy effects.
@@ 2.1.1 Sommer parameter $r_{0}$ and setting the scale
can be used to define a length scale $r_{0}$, the so-called Sommer parameter [24], as follows
$r^{2}\frac{\mathrm{d}}{\mathrm{d}r}V(r)|_{r=r_{0}}=c$ (9)
in units of the spatial lattice spacing $a_{s}$. In QCD, the physical value of $r_{0}$ is known to be around 0.5fm0.5\ $\mathrm${fm} for $c=1.65$ , but, in the U(1)(1) theory at hand its physical value is unknown. However, this is not relevant for our procedure, as we only need an observable in units of the spatial lattice spacing with a well defined continuum limit. We use a value of $c=1.65$, because it turns out to be in the linear region of the potential and we stick to the notation $r_{0}/a_{s}$. Once the potentials are parametrised using the form eq. 8, the corresponding value of $r_{0}/a_{s}$ can be determined. It allows one to fix the spatial lattice spacing: if in two simulations with parameters $(\beta,\xi)$ and $(\beta^{\prime},\xi^{\prime})$ the same $r_{0}/a_{s}$ is measured within uncertainties, both simulations exhibit the same $a_{s}$.
@@ 2.2 Hamiltonian
The Kogut-Susskind Hamiltonian for the pure U(1) lattice gauge theory in (2+1) dimensions is given by [2]
@@
where $\vec{A}_{\vec{r},\mu}$ is the discretized vector field in the compact formulation, i.e., the values of $ag\vec{A}_{\vec{r},\mu}$ are constrained to lie within the interval $0,2\pi)$. Note that $\hat{U}_{\vec{r},\mu}$ is a unitary operator. The commutation relations between the electric field operator $\hat{E}_{\vec{r},\nu}$ and the link operator $\hat{U}_{\vec{r^{\prime}},\mu}$ read
@@
[…]
where $Q_{\vec{r}}$ are the static charges. Instead of considering the full Hilbert space and enforcing Gauss’s law, we formulate the theory directly on the gauge-invariant subspace, by using the Gauss’s law constraints to eliminate certain degrees of freedom [25, 26, 27, 28, 29].
For a numerical implementation of the Hamiltonian, the gauge degrees of freedom must be truncated to a finite dimension because the electric field values are unbounded, resulting in an infinite-dimensional Hilbert space for these degrees of freedom. The continuous $U(1)$ gauge group can be discretized in the electric basis to the group of integers $\mathbb{Z}_{2l+1}$. The integer $l$ sets the truncation level, i.e., the discretized gauge fields are constrained to integer values within the range [l,l][-l,l] [25]. The total Hilbert space dimension is $(2l+1)^{N}$, where $N$ is the number of gauge fields.
@@
[…]
When discretizing a gauge theory, one needs to give up either unitarity or the exact commutation relations between the electric field and link operators in eq. 13 and eq. 14. In our case, the commutation relations are preserved for the truncated operators, but unitarity is lost, $\hat{U}^{{\dagger}}_{\vec{r},\mu}\hat{U}_{\vec{r},\mu}\neq\mathbbm{1}$. This can be seen from the matrix representation of the link operators [27],
@@
However, unitarity is recovered in the limit of $l\to\infty$. The resulting errors due to the finite truncation parameter $l$ have been investigated in Refs. [30, 31]. Alternative methods for defining the electric field and link operators have been explored in Refs. [32, 33, 34, 35].
In order to analytically derive the Kogut-Susskind Hamiltonian in eq. 11 from the Wilson action in eq. 1, one needs to employ the transfer matrix method [3]. This derivation has been performed for various quantum field theories in arbitrary dimensions, for example, for studying transport coefficients [36] and the topological $\theta$-term of (non)Abelian lattice gauge theories in (3+1) dimensions [28]. Since the Wilson action is defined on a $(d+1)$-dimensional space-time lattice and the Kogut-Susskind Hamiltonian is defined on a $d$-dimensional spatial lattice, the limit of $a_{t}\to 0$ has to be taken when deriving eq. 11 from eq. 1 using the transfer matrix formalism [3].
@@ 3.1 Lagrangian
Details of the algorithm can be found in Ref. [37] (see in particular the arXiv version).
We provide our codes for the generation of the gauge ensembles and for the analysis of the ensembles in [38] [^1] Please see the README-files for more information on how to generate the ensembles and how to replicate the analysis.. For the analysis, we have used the library [39].
@@ 3.2 Hamiltonian
As discussed in section 2.2, we use the Gauss’s law to eliminate some gauge degrees of freedom, thereby restricting the theory to the gauge-invariant space. More specifically, we treat the Gauss’s law in eq. 15 as a set of constraints on the electric operators, and solve this set of equations over the electric operators. While there are $N$ Gauss’s law constraints, they are not independent, since there is a conservation of charges, which, in the pure gauge case, means the constraints sum to zero. Therefore, there are only N1N-1 independent constraints, which allows us to express N1N-1 arbitrary electric field operators, i.e., effectively eliminating them, in terms of the remaining ones. Since the eliminated electric fields do not contribute directly to the dynamics, their corresponding link operators become identities. For a two-dimensional $L\times L$ square lattice, where $N=L^{2}$, $L^{2}-1$ out of the $2L^{2}$ gauge degrees of freedom are eliminated. Thus, the Hamiltonian is expressed in terms of the $L^{2}+1$ remaining gauge fields. This reduces the number of basis states from $(2l+1)^{2L^{2}}$ to $(2l+1)^{L^{2}+1}$, which in practice, alleviates the computational costs significantly.
@@
where $V$ is the number of plaquettes in the lattice. Note that we set the lattice spacing $a=1$ throughout our simulations. Truncating the electric field on the dynamical links to a range of [l,l][-l,l] can in principle lead to configurations where the links that have been eliminated implicitly carry an electric flux exceeding this range. These would violate discrete symmetries of the Hamiltonian and are unphysical. To ensure that our results are not affected by such effects, the simulations are carried out over a range of $1/g^{2}$ in $(0,10]$ and for $l\in\{1,2,3,4\}$. We find that the values of $\langle\mathcal{P}\rangle$ obtained for $l=2,3,4$ agree with each other up to $1/g^{2}=1.5$, and are thus not affected by truncation artifacts. Beyond these values of the inverse coupling, the values start to deviate for different $l$ values, indicating that the simulations are no longer reliable.